OpenStudy (anonymous):

Let f(x) be mx-1 + 1/x. Then the smallest value of the constant m such that f(x)>=0 for every x>0.

OpenStudy (anonymous):

may be or maybe not....

OpenStudy (anonymous):

try solving the equation f(x)=0 for x

OpenStudy (anonymous):

constraint is there for x so cant

OpenStudy (anonymous):

ignore the constraint for now

OpenStudy (anonymous):

m x-1+1/x=0. multiply both sides by x

OpenStudy (anonymous):

nope that will not work

OpenStudy (anonymous):

once you solve for x in terms of m, all you need to find is a value of m for which the x values are almost imaginary. this occurs when the discriminant is zero.

OpenStudy (anonymous):

it gives an open interval at that value, 'that value' is the answer...

OpenStudy (anonymous):

so any idea bout calculus?

OpenStudy (anonymous):

x={(1+sqrt(1-4m))/2m,(1-sqrt(1-4m))/2m}. for the discriminant to be zero, m must equal 1/4.

OpenStudy (anonymous):

its in an open interval so cant say 1/4

OpenStudy (anonymous):

you mean the interval for the x values?

OpenStudy (anonymous):

what u are saying is x>1/4 so it will not include 1/4

OpenStudy (anonymous):

for m>1/4, f(x)>0. thus, for m=1/4, f(x)>=0.

OpenStudy (anonymous):

nope....cant say like that

OpenStudy (anonymous):

ok sure. using calculus, set the first derivative equal to zero

OpenStudy (anonymous):

to find the minimum point of the function

OpenStudy (anonymous):

you get x=1/sqrt(m). now plug this value back into the original equation, and set f(1/sqrt(m)) equal to zero.

OpenStudy (anonymous):

d/dx (m x - 1 + 1/x)=0 m-1/x^2=0 1/x^2=m x=1/sqrt(m)

OpenStudy (anonymous):

f(1/sqrt(m))=m/sqrt(m)-1+sqrt(m)=0. m-sqrt(m)+m=0 2m-sqrt(m)=0 sqrt(m)=2m m=1/4

OpenStudy (anonymous):

thanks buddy...

OpenStudy (anonymous):

:)